# A Single-Loop First-Order Algorithm for Linearly Constrained Bilevel Optimization

### v2

@@ 1 Introduction
In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (pedregosa2016hyperparameter; franceschi2018bilevel; mackay2019self), data hypercleaning (shaban2019truncated), meta learning (rajeswaran2019meta; ji2020convergence), reinforcement learning (sutton2018reinforcement; hong2023two) and neural architecture search (liu2018darts; liang2019darts+); see survey papers (zhang2024introduction; liu2021investigating; sinha2017review) for additional discussions. While numerous works for unconstrained BLO problems have been proposed (ghadimi2018approximation; liu2020generic; ji2021bilevel; dagreou2022framework; hong2023two; kwon2023fully), studies focusing on constrained BLO problems are relatively limited.
@@
For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$. This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (yang2022decentralized), hyperparameter optimization of constrained learning problems (xu2023efficient) and adversarial training (zhang2022revisiting) and draw significant attentions (tsaknakis2022implicit; khanduri2023linearly; kornowski2024first).
A popular approach for solving unconstrained BLO is implicit gradient descent (ghadimi2018approximation; ji2021bilevel; ji2022will; chen2022single; khanduri2021near). For constrained BLO, several studies have extended this approach to accommodate different constraint settings (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating). However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating). The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
Recently, some first-order methods (kwon2023penalty; yao2024constrained; yao2024overcoming; jiang2024primal; kornowski2024first) have been proposed for addressing constrained BLO problems. Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem. For example, yao2024constrained; yao2024overcoming reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem. However, neither yao2024constrained nor yao2024overcoming establishes clear relationships between the stationary points of their approximated problems and the original one.
Works most closely related to ours are those by kwon2023penalty and jiang2024primal, both of which reformulated the problem (1) as
@@
where $\Phi_{\delta}(x,y,z)=f(x,y)+\frac{1}{\delta}[g(x,y)-g(x,z)]$. Based on this reformulation, kwon2023penalty proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and jiang2024primal proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. However, kwon2023penalty only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration. Algorithms in jiang2024primal require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation. Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (kwon2023penalty; jiang2024primal) for coupled constraints $\mathcal{Y}(x)$.
@@
  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem, and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies). The simple single-loop structure also makes our algorithm easier to implement in practice compared to jiang2024primal.
@@
Table 1: Comparison of our paper with (kwon2023penalty; jiang2024primal). More detailed introductions and discussions of other related works can be found in Section˜2. Here, the “Complexity” means the iteration complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$ (3).
@@
kwon2023penalty $y\in\mathcal{Y}$, $\mathcal{Y}$ is a convex and compact set $O(\epsilon^{-3}\log(\epsilon^{-1}))$ single/double
jiang2024primal $h(x,y)\leq 0$, LICQ holds $O(\epsilon^{-5}\log(\epsilon^{-1}))$ triple
jiang2024primal $B(x)+A(x)y\leq 0$, $A(x)$ is full row rank $O(\epsilon^{-3}\log(\epsilon^{-1}))$ double
@@ 2 Related works
BLO without constraints. One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (pedregosa2016hyperparameter). It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (ghadimi2018approximation). Later works improved the convergence rates and studied the gradient descent methods under various settings (ji2021bilevel; ji2022will; chen2022single; khanduri2021near; yang2023accelerating). Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (maclaurin2015gradient; grazzi2020iteration; liu2021towards; bolte2022automatic). Recently, penalty-based methods have gained traction as a promising approach for solving BLO. Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (liu2021value; mehra2021penalty; liu2022bome; kwon2023fully; shen2023penalty; gao2022value; lu2024slm; lu2024first).
@@
BLO with constraints. There are two primary types of methods for solving constrained bilevel optimization problems. One is based on the implicit gradient method. Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (khanduri2023linearly). tsaknakis2022implicit proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm. khanduri2023linearly proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$. xu2023efficient used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$. However, they only provided an asymptotic convergence analysis of their algorithm. xiao2023alternating proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem. However, these implicit gradient-type algorithms (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; xiao2023alternating) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
@@
Another commonly used approach for solving constrained BLO problems is based on penalty reformulation. For example, lu2024first reformulated unconstrained and constrained BLO problems as structured minimax problems and introduced first-order methods with guarantees for finding $\epsilon$-KKT solutions. yao2024constrained reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function. However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly. yao2024overcoming reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm. Compared to yao2024constrained, yao2024overcoming did not need the projection operator to the coupled constraint set. However, both yao2024constrained and yao2024overcoming did not establish very clear relationships between the stationary points of their approximated problems and the original one. For example, yao2024overcoming only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one. Recently, tsaknakis2023implicit; jiang2024barrier proposed algorithms based on barrier approximation approach for constrained BLO problems. However, their algorithms also require the computations for the Hessian matrix.
@@
kwon2023penalty and jiang2024primal considered the same reformulation as ours. kwon2023penalty studied the case where the LL variables $y\in\mathcal{Y}$ are independent of $x$ and characterized the conditions under which the values and derivatives of $\Phi$ and $\Phi_{\delta}$ can be $O(\delta)$-close for $y\in\mathcal{Y}$ constraints. Compared with kwon2023penalty, we prove similar results under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. Moreover, the algorithms in kwon2023penalty require the implementation of the projection operator to $\mathcal{Y}$ at each iteration, which can be costly. jiang2024primal studied the coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. While jiang2024primal considered more general constraints than ours, however, it did not characterize the gap between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$. Our Theorem˜4.9 provides further justifications for their reformulation in coupled constraints. Moreover, compared with the double- and triple-loop algorithms in kwon2023penalty, we propose a single-loop algorithm SFLCB and prove an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$.
@@
Recently, kornowski2024first also proposed first-order methods for linearly constrained BLO. Especially, they proved a nearly optimal convergence rate $\tilde{O}(\epsilon^{-2})$ for linear equality constraints and proposed algorithms that can attain $(\delta,\epsilon)$-Goldstein stationarity for linear inequality constraints. However, their convergence rates for linear inequality constraints either have additional dependence on dimension $d$ (such as $\tilde{O}(d\delta^{-1}\epsilon^{-3})$) or need additional assumptions to access the exact optimal dual variable (such as $\tilde{O}(\delta^{-1}\epsilon^{-4})$), while we do not require the exact optimal dual variable assumption. Compared with the double-loop algorithms in kornowski2024first, our proposed single-loop one is easier to implement in practice. Moreover, our techniques are also different from theirs under linear inequality constraints, thereby highlighting the distinct contributions and independent interests of our work.
@@ 4 Reformulation
Similar to Theorem 3.8 in kwon2023penalty, we have the following theorem to bound the difference between $\Phi$ and $\Phi_{\delta}$, as well as $y^{*}(x)$ and $y_{\delta}^{*}(x)$ in the coupled constraints.
@@
Then, similar to Lemmas 2 and 3 in jiang2024primal, we have the following lemma.
@@
While the gradients of $\Phi_{\delta}(x)$ exists under LICQ conditions, for general problem (1), $\Phi(x)$ is not guaranteed to be differentiable. For example, khanduri2023linearly provides an example where the LICQ condition holds, $\Phi(x)$ is non-differentiable at some points. However, if a given $x$ satisfies the following conditions, then $\nabla\Phi(x)$ exists at $x$.
@@
Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (tsaknakis2022implicit; khanduri2023linearly; xu2023efficient; jiang2024barrier; kwon2023penalty) to ensure the existence of $\nabla\Phi(x)$.
Lemma 4.7 (Theorem 2 in xu2023efficient).
@@
˜4.8 is made for $t\in[0,\delta]$. When $\delta$ is sufficiently small, i.e., $\delta=O(\epsilon)$, $y^{*}(x)$ and $y^{*}_{t}(x)$ are very close according to Theorem˜4.1. Thus, we expect that for $t\in[0,\delta]$, $y^{*}_{t}(x)$ will have similar properties as $y^{*}(x)$. Similar assumptions have also been used in kwon2023penalty to establish the non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$.
@@
Similar non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$ has been established in kwon2023penalty; however, their bound is established only for the LL constraints $\mathcal{Y}$ that do not depend on $x$. Our Theorem˜4.9 provides a more general theoretical justification for the validity of the reformulation (3) for coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.

### v3

@@ 1 Introduction
In recent years, bilevel optimization (BLO) has gained significant popularity for addressing a wide range of modern machine learning problems, such as hyperparameter optimization (Pedregosa, 2016; Franceschi et al., 2018; MacKay et al., 2019), data hypercleaning (Shaban et al., 2019), meta learning (Rajeswaran et al., 2019; Ji et al., 2020), reinforcement learning (Sutton, 2018; Hong et al., 2023) and neural architecture search (Liu et al., 2018; Liang et al., 2019); see survey papers (Zhang et al., 2024; Liu et al., 2021a; Sinha et al., 2017) for additional discussions. While numerous works for unconstrained BLO problems have been proposed (Ghadimi and Wang, 2018; Liu et al., 2020; Ji et al., 2021; Dagréou et al., 2022; Hong et al., 2023; Kwon et al., 2023a), studies focusing on constrained BLO problems are relatively limited.
@@
For this setting, we develop a single loop algorithm for the special case h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, where $B\in\mathbb{R}^{d_{h}\times d_{x}}$ and $A\in\mathbb{R}^{d_{h}\times d_{y}}$. This special class of constrained BLO problems covers a wide class of applications, including distributed optimization (Yang et al., 2022), hyperparameter optimization of constrained learning problems (Xu and Zhu, 2023) and adversarial training (Zhang et al., 2022) and draw significant attentions (Tsaknakis et al., 2022; Khanduri et al., 2023; Kornowski et al., 2024).
A popular approach for solving unconstrained BLO is implicit gradient descent (Ghadimi and Wang, 2018; Ji et al., 2021, 2022; Chen et al., 2022; Khanduri et al., 2021). For constrained BLO, several studies have extended this approach to accommodate different constraint settings (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Xiao et al., 2023). However, these implicit gradient-based methods in constrained BLO necessitate computing the Hessian matrix of the lower-level problem (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Xiao et al., 2023). The potential computational challenges associated with the Hessian matrix limit their practical applicability for large-scale problems.
Recently, some first-order methods (Kwon et al., 2023b; Yao et al., 2024b, a; Jiang et al., 2024a; Kornowski et al., 2024) have been proposed for addressing constrained BLO problems. Most of those works considered transforming the original problem (1) into a single-level one and trying to find the stationary points of the reformulated problem. For example, Yao et al. (2024b, a) reformulated the original problem into some approximated functions and proposed single-loop algorithms for finding the stationary point of the approximated problem. However, neither Yao et al. (2024b) nor Yao et al. (2024a) establishes clear relationships between the stationary points of their approximated problems and the original one.
Works most closely related to ours are those by Kwon et al. (2023b) and Jiang et al. (2024a), both of which reformulated the problem (1) as
@@
where $\Phi_{\delta}(x,y,z)=f(x,y)+\frac{1}{\delta}[g(x,y)-g(x,z)]$. Based on this reformulation, Kwon et al. (2023b) proposed algorithms for solving BLO with LL constraints $y\in\mathcal{Y}$ and Jiang et al. (2024a) proposed algorithms for solving coupled LL constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. However, Kwon et al. (2023b) only considered the LL constraints $\mathcal{Y}$ that are independent of $x$ and their methods require projection oracle to $\mathcal{Y}$ at each iteration. Algorithms in Jiang et al. (2024a) require complex double or triple loops, resulting in sub-optimal convergence rates and difficult implementation. Moreover, the connection between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$ is not discussed in (Kwon et al., 2023b; Jiang et al., 2024a) for coupled constraints $\mathcal{Y}(x)$.
@@
  • Based on this reformulation and equipped with augmented Lagrangian methods, we proposed SFLCB, a single-loop, first-order algorithm for linearly constrained bilevel optimization problem, and provide rigorous analyses of its non-asymptotic convergence rates, achieving an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$ compared to prior works (See Table˜1 for a more comprehensive comparison of our work with previous studies). The simple single-loop structure also makes our algorithm easier to implement in practice compared to Jiang et al. (2024a).
@@
Table 1: Comparison of our paper with (Kwon et al., 2023b; Jiang et al., 2024a). More detailed introductions and discussions of other related works can be found in Section˜2. Here, the “Complexity” means the iteration complexity needed to achieve the $\epsilon$-stationary point of $\Phi_{\delta}$ (3).
@@
Kwon et al. (2023b) $y\in\mathcal{Y}$, $\mathcal{Y}$ is a convex and compact set $O(\epsilon^{-3}\log(\epsilon^{-1}))$ single/double
Jiang et al. (2024a) $h(x,y)\leq 0$, LICQ holds $O(\epsilon^{-5}\log(\epsilon^{-1}))$ triple
Jiang et al. (2024a) $B(x)+A(x)y\leq 0$, $A(x)$ is full row rank $O(\epsilon^{-3}\log(\epsilon^{-1}))$ double
@@ 2 Related works
BLO without constraints. One popular approach for solving unconstrained BLO is to use implicit gradient descent methods (Pedregosa, 2016). It is well established that when the LL problem is strongly convex and unconstrained, $y^{*}(x)=\mathop{\rm argmin}_{y}g(x,y)$ exists and is differentiable, and the gradient of the hyper-objective can be calculated by $\nabla\Phi(x)=\nabla_{x}f(x,y)+(\nabla y^{*}(x))^{\top}\nabla_{y}f(x,y^{*}(x))$ (Ghadimi and Wang, 2018). Later works improved the convergence rates and studied the gradient descent methods under various settings (Ji et al., 2021, 2022; Chen et al., 2022; Khanduri et al., 2021; Yang et al., 2023). Another popular approach is based on iterative differentiation, which iteratively solves the LL problems and computes $\nabla y^{*}(x)$ to approximate the hypergradient (Maclaurin et al., 2015; Grazzi et al., 2020; Liu et al., 2021c; Bolte et al., 2022). Recently, penalty-based methods have gained traction as a promising approach for solving BLO. Those works usually reformulate the original BLO as the single-level one and use the first-order methods to find the stationary point of the reformulated problems (Liu et al., 2021b; Mehra and Hamm, 2021; Liu et al., 2022; Kwon et al., 2023a; Shen and Chen, 2023; Gao et al., 2022; Lu, 2024; Lu and Mei, 2024).
@@
BLO with constraints. There are two primary types of methods for solving constrained bilevel optimization problems. One is based on the implicit gradient method. Generally, when the LL problem has constraints, the differentiabilities of $y^{*}(x)$ and $\Phi(x)$ are not guaranteed (Khanduri et al., 2023). Tsaknakis et al. (2022) proved the existence of $\nabla\Phi(x)$ under additional assumptions for linearly constraint $Ay\leq b$ and proposed an implicit gradient-type double-loop algorithm. Khanduri et al. (2023) proposed a perturbation-based smoothing technique to compute the approximate implicit gradient for linearly constraint $Ay\leq b$. Xu and Zhu (2023) used Clarke subdifferential to approximate the non-differentiable implicit function $\Phi$. However, they only provided an asymptotic convergence analysis of their algorithm. Xiao et al. (2023) proved the existence of $\nabla\phi$ where the LL has equality constraints $Ay+H(x)=c$, and introduced an alternating projected SGD approach to solve this problem. However, these implicit gradient-type algorithms (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Xiao et al., 2023) require the computations for the Hessian matrix of the LL problems, which potentially limit their practical applicability for large-scale problems.
@@
Another commonly used approach for solving constrained BLO problems is based on penalty reformulation. For example, Lu and Mei (2024) reformulated unconstrained and constrained BLO problems as structured minimax problems and introduced first-order methods with guarantees for finding $\epsilon$-KKT solutions. Yao et al. (2024b) reformulated the original problem into a proximal Lagrangian value function and proposed a single-loop, first-order method to find the stationary points of the reformulated value function. However, their algorithm requires the implementation of the projection operator on $\mathcal{C}=\{x,y|h(x,y)\leq 0\}$ at each iteration, which can be potentially costly. Yao et al. (2024a) reformulated the original problem into a doubly regularized gap function and proposed a single-loop, first-order algorithm. Compared to Yao et al. (2024b), Yao et al. (2024a) did not need the projection operator to the coupled constraint set. However, both Yao et al. (2024b) and Yao et al. (2024a) did not establish very clear relationships between the stationary points of their approximated problems and the original one. For example, Yao et al. (2024a) only provided an asymptotic relationship between the original problem and their reformulated one, i.e., as their penalty parameter approaches infinite, their reformulated problem is equivalent to the original one. Recently, Tsaknakis et al. (2023); Jiang et al. (2024b) proposed algorithms based on barrier approximation approach for constrained BLO problems. However, their algorithms also require the computations for the Hessian matrix.
@@
Kwon et al. (2023b) and Jiang et al. (2024a) considered the same reformulation as ours. Kwon et al. (2023b) studied the case where the LL variables $y\in\mathcal{Y}$ are independent of $x$ and characterized the conditions under which the values and derivatives of $\Phi$ and $\Phi_{\delta}$ can be $O(\delta)$-close for $y\in\mathcal{Y}$ constraints. Compared with Kwon et al. (2023b), we prove similar results under coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. Moreover, the algorithms in Kwon et al. (2023b) require the implementation of the projection operator to $\mathcal{Y}$ at each iteration, which can be costly. Jiang et al. (2024a) studied the coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$. While Jiang et al. (2024a) considered more general constraints than ours, however, it did not characterize the gap between the stationary point of the reformulated function $\Phi_{\delta}$ and the original hyper-objective $\Phi$. Our Theorem˜4.9 provides further justifications for their reformulation in coupled constraints. Moreover, compared with the double- and triple-loop algorithms in Kwon et al. (2023b), we propose a single-loop algorithm SFLCB and prove an improvement in the convergence rate from $O(\epsilon^{-3}\log(\epsilon^{-1}))$ to $O(\epsilon^{-3})$.
@@
Recently, Kornowski et al. (2024) also proposed first-order methods for linearly constrained BLO. Especially, they proved a nearly optimal convergence rate $\tilde{O}(\epsilon^{-2})$ for linear equality constraints and proposed algorithms that can attain $(\delta,\epsilon)$-Goldstein stationarity for linear inequality constraints. However, their convergence rates for linear inequality constraints either have additional dependence on dimension $d$ (such as $\tilde{O}(d\delta^{-1}\epsilon^{-3})$) or need additional assumptions to access the exact optimal dual variable (such as $\tilde{O}(\delta^{-1}\epsilon^{-4})$), while we do not require the exact optimal dual variable assumption. Compared with the double-loop algorithms in Kornowski et al. (2024), our proposed single-loop one is easier to implement in practice. Moreover, our techniques are also different from theirs under linear inequality constraints, thereby highlighting the distinct contributions and independent interests of our work.
@@ 4 Reformulation
Similar to Theorem 3.8 in Kwon et al. (2023b), we have the following theorem to bound the difference between $\Phi$ and $\Phi_{\delta}$, as well as $y^{*}(x)$ and $y_{\delta}^{*}(x)$ in the coupled constraints.
@@
Then, similar to Lemmas 2 and 3 in Jiang et al. (2024a), we have the following lemma.
@@
While the gradients of $\Phi_{\delta}(x)$ exists under LICQ conditions, for general problem (1), $\Phi(x)$ is not guaranteed to be differentiable. For example, Khanduri et al. (2023) provides an example where the LICQ condition holds, $\Phi(x)$ is non-differentiable at some points. However, if a given $x$ satisfies the following conditions, then $\nabla\Phi(x)$ exists at $x$.
@@
Note that ˜4.5, 4.6 are commonly used in constrained BLO literature (Tsaknakis et al., 2022; Khanduri et al., 2023; Xu and Zhu, 2023; Jiang et al., 2024b; Kwon et al., 2023b) to ensure the existence of $\nabla\Phi(x)$.
Lemma 4.7 (Theorem 2 in Xu and Zhu (2023)).
@@
˜4.8 is made for $t\in[0,\delta]$. When $\delta$ is sufficiently small, i.e., $\delta=O(\epsilon)$, $y^{*}(x)$ and $y^{*}_{t}(x)$ are very close according to Theorem˜4.1. Thus, we expect that for $t\in[0,\delta]$, $y^{*}_{t}(x)$ will have similar properties as $y^{*}(x)$. Similar assumptions have also been used in Kwon et al. (2023b) to establish the non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$.
@@
Similar non-asymptotic bound for $\|\nabla\Phi(x)-\nabla\Phi_{\delta}(x)\|$ has been established in Kwon et al. (2023b); however, their bound is established only for the LL constraints $\mathcal{Y}$ that do not depend on $x$. Our Theorem˜4.9 provides a more general theoretical justification for the validity of the reformulation (3) for coupled constraints $\mathcal{Y}(x)=\{y\in\mathbb{R}^{d_{y}}|h(x,y)\leq 0\}$.